### v1

@@ Abstract
For a subfield $\mathbb{K}$ of the field $\mathbb{C}$ of complex numbers, we consider curve and divisorial valuations on the algebra $\mathbb{K}[[x,y]]$ of formal power series in two variables with the coeficients in $\mathbb{K}$. We compute the semigroup Poincaré series and the classical Poincaré series of a valuation.
@@ 1 Introduction
where $J^{\mathbb{C}}(v):=\{f\in{\mathcal{O}}_{\mathbb{C}^{2},0}:\nu(f)\geq v\}$. The Poincaré series of one curve valuation was known to specialists for many years. Computations can be found, e. g., in [4]. The Poincaré series of one divisorial valuation was computed in [9]. The Poincaré series of a collection of curve valuations was computed in [5]. It was shown that this series coincides with the Alexander polynomial in several variables of the corresponding algebraic link. This implies that the Poincaré series determines the topology of the union of the corresponding curves.
@@ 2 Curve valuations on functions over a field
Let $\mathbb{K}$ be a subfield of the field $\mathbb{C}$ of complex numbers and let ${\mathcal{E}}_{\mathbb{K}^{2},0}$ be the algebra of germs of holomorphic functions in two variables with the Taylor coefficients from $\mathbb{K}$. A (discrete, rank one) valuation on ${\mathcal{E}}_{\mathbb{K}^{2},0}$ (defined as above) is the restriction of a valuation on ${\mathcal{O}}_{\mathbb{C}^{2},0}$: see, e. g., [10].
In the framework of the discussion below, without changes it is possible to consider, instead of the algebra ${\mathcal{E}}_{\mathbb{K}^{2},0}$, the algebra $\mathbb{K}[[x,y]]$ of formal power series in two variables. We shall refer to ${\mathcal{E}}_{\mathbb{K}^{2},0}$ because of some convenience of the language.
Let $(C,0)$ be a plane algebroid branch on $(\mathbb{C}^{2},0)$, that is $C$ (after an appropriate change of the coordinates in $\mathbb{C}^{2}$) is defined by
$x=x(\tau)=\tau^{m},\\ y=y(\tau)=\sum_{i\geq m}c_{i}\tau^{i}\in\mathbb{C}[[\tau]]\,.$ (2)
(We assume that the parametrization (2) is irreducible, i. e., the greatest common divisor of $m$ and of all $i$ such that $c_{i}\neq 0$ is equal to $1$.) The branch $C$ defines a valuation $\nu=\nu_{C}$ on the algebra ${\mathcal{O}}_{\mathbb{C}^{2},0}$ of holomorphic function germs in $x$ and $y$. For $f\in{\mathcal{O}}_{\mathbb{C}^{2},0}$, $\nu(f)$ is the degree of the leading term in the series $f(x(\tau),y(\tau))\in\mathbb{C}[[\tau]]$:
@@
The $\mathbb{K}$-Poincaré series of the branch $C$ is
$P_{C}(t)(=P_{C}^{\mathbb{K}}(t))=\sum_{v=0}^{\infty}\dim_{\mathbb{K}}\left(J^{\mathbb{K}}(v)/J^{\mathbb{K}}(v+1)\right)\cdot t^{v}.$ (4)
Example 1.
If $\mathbb{K}=\mathbb{R}$, $P_{C}(t)$ is the classical Poincaré series of the valuation $\nu$ computed in [7].
@@
If $\mathbb{K}=\mathbb{C}$, all the coefficients of the series (4) are equal to 00 or $1$. This is not the case, in general, for a proper subfield $\mathbb{K}\subset\mathbb{C}$. In particular, if $\mathbb{K}=\mathbb{R}$ (the field of real numbers), all the coefficients are equal to 00, $1$, or $2$ ([7]).
Let $S^{\mathbb{K}}_{C}:=\{\nu(f):f\in{\mathcal{E}}_{\mathbb{K}^{2},0}\}\subset\mathbb{Z}_{\geq 0}$ be the semigroup of values of the valuation $\nu_{C}$ on ${\mathcal{E}}_{\mathbb{K}^{2},0}$.
@@
Here we compute the $\mathbb{K}$-semigroup Poincaré series and the $\mathbb{K}$-Poincaré series for one curve or divisorial valuation.
@@ 3 The resolution process of a branch
Let $G$ be the Galois group of the extention $\mathbb{C}/\mathbb{K}$. i. e. the group of automorphisms of $\mathbb{C}$ which are trivial on $\mathbb{K}$. Let us consider the minimal $G$-invariant resolution process of the branch $C$ (by a sequence of blow-ups). This means the following. When we blow-up the origin in $\mathbb{C}^{2}$, we have got one of the following three possibilities.
@@
    There are finitely many different branches in the $G$-orbit of the strict transform $\widetilde{C}$ of the branch $C$, all of them are smooth and they intersect the exceptional divisor transversally at pairwise different points. We get (the minimal embedded $G$-invariant) resolution of $C$ and the process is finished.
@@
    The branches from the $G$-orbit of $\widetilde{C}$ (whose number may be infinite) intersect the exceptional divisor at finite number of points, but either they are not smooth, or do not intersect the exceptional divisor transversally, or (some) different branches from the orbit intersect it at the same point. In this case one has to continue the process.
@@
    The branches from the $G$-orbit of $\widetilde{C}$ intersect the exceptional divisor at infinitely many different points. In this case the process is terminated.
In Case 2 one has to blow-up (all) the intersection points of the branches from the $G$-orbit of $\widetilde{C}$ with the exceptional divisor (their number being finite). One gets one of the analogues the cases 1–3 again. Etc.
As the result we arrive to one of the following three situations.
  1. I
    There are finitely many branches in the $G$-orbit of $C$ and they are resolved at a certain step. The dual graph of the resolution looks like in Figure 1.
@@
The Galois group $G$ acts on the graph $\Gamma$ and the quotient $\check{\Gamma}$ by the action looks like in Figure 2 (and further on).
@@
The graph $\Gamma$ contains a subgraph which geometrically coincides with the dual graph of a usual resolution of the curve $C$: the lower part of $\Gamma$ on Figure 1. (We do not consider this resolution below.) We use “short” notations (by greek letters) for verticies on this subgraph (and also for the corresponding verticies of the graph $\check{\Gamma}$). All other verticies are obtain from them by the action of elements of the Galois group $G$.
@@
In Figure 1, $\sigma_{i}$, $i=0,\ldots,g$, (and also their images under the action of $G$) are the dead ends of the graph ($g$ is the number of the Puiseux pairs of the curve $C$), $\tau_{i}$, $i=1,\ldots,g$, (and also their images) are the rupture points of it, $\rho_{j}$, $j=1,\ldots,s$, ($s$ is equal to $\infty$ in Case II) are the splitting points corresponding to the changes of the isotropy subgroups for the $G$-action on $\Gamma$ (that is the splitting points for usual resolution of different branches from the orbit of $C$, there are finitely many of them in Case I infinitely many in Case II.
Figure 2 shows $\check{\Gamma}$ in the case when the process of the $\mathbb{K}$-resolution of $C$ is not finish when $C$ itself is resolved. Otherwise $\delta_{C}$ coincides with $\tau_{g}$.
@@
Figure 2: The resolution graph $\check{\Gamma}$ in Case I (with arbitrary $g$ and $s$).
  1. II
@@
    The described process goes on without termination. This means that we get an infinite sequence of blow-ups. The corresponding dual graph of the process looks like in Figure 3. One can say that $s$ is equal to infinity in this case.
@@
    Figure 3: The resolution graph $\check{\Gamma}$ in Case II (with $s=\infty$).
  2. III
    The process is terminated at a certain step. This means that at this step the branches from the $G$-orbit of $\widetilde{C}$ intersect each of the new-born components of the exceptional divisor at infinitely many points. The dual graph of the process looks like in Figure 4.
@@ 2. III
    In the latter case, the valuation coincides with $n\nu$, where $\nu=\nu_{\rho_{s}}$ is the divisorial valuation defined by $E_{\rho_{s}}$ and $n$ is the intersection multiplicity $[E_{\rho_{s}},\widetilde{C}]$ of the strict transform $\widetilde{C}$ of the curve $C$ with the component $E_{\rho_{s}}$.
    Figure 4: The resolution graph $\check{\Gamma}$ in Case III.
Cases I–III take place in the following situations.
  1. I
    All the coefficients in the series (2) are from a finite extension of $\mathbb{K}$. This is equivalent to the condition that the orbit $GC$ has finitely many elements.
  2. II
  3. III
@@
Let $\mathbb{K}\subset\mathbb{C}$ (e. g., $\mathbb{K}=\mathbb{Q}$) and let $F\in\mathbb{K}[[x,y]]$ be irreducible. The local ring of the irreducible curve (over $\mathbb{K}$) defined by $F$ is $R=\mathbb{K}[[x,y]]/(F)$. The integral closure $\overline{R}$ in its quotient field is the valuation ring $R_{\omega}$ of a discrete valuation $\omega$, its residual field $\mathbb{K}^{\prime}$ is a finite algebraic extension of $\mathbb{K}$ and could be realized and therefore identified with a subfield of $\mathbb{C}$. The equation $F$ defines a reduced algebroid curve over the complex field $\mathbb{C}$, each branch corresponds to an irreducible factor of $F$ in $\mathbb{C}[[x,y]]$. Let $C$ be one of these branches. The valuation $\nu=\nu_{C}$ defined by $C$ on $\mathbb{K}[[x,y]]$ coincides with the valuation $\omega$ (seen as a valuation on $\mathbb{K}[[x,y]]$ forgeting the quotient by $F$) and the Poincaré series (over $\mathbb{K}$) defined by $\nu$ is the Poincaré series defined by $\omega$. Note that all the branches are equisingular, the minimal resolution of $\{F=0\}$ (in the complex setting) coincides with the $\mathbb{K}$-resolution of $\nu_{C}$ and all the valuations $\nu_{C}$, for complex branches $C$ of $\{F=0\}$, restrict to the same valuation $\omega$ in $\mathbb{K}[[x,y]]$. Moreover, if $G^{\prime}$ is the Galois group of the extension $\mathbb{K}\subset{\mathbb{L}}$, ${\mathbb{L}}$ the normal closure of $\mathbb{K}^{\prime}$ in $\mathbb{C}$, $G^{\prime}$ is the quotient of $G$ by the isotropy group of the field $\mathbb{L}$. This situation just corresponds to the case I described above.

### v2

@@ Abstract
For a subfield $\mathbb{K}$ of the field $\mathbb{C}$ of complex numbers, we consider curve and divisorial valuations on the algebra $\mathbb{K}[[x,y]]$ of formal power series in two variables with the coefficients in $\mathbb{K}$. We compute the semigroup Poincaré series and the classical Poincaré series of a (discrete, rank one) valuation on it.
@@ 1 Introduction
where $J^{\mathbb{C}}(v):=\{f\in{\mathcal{O}}_{\mathbb{C}^{2},0}:\nu(f)\geq v\}$. The Poincaré series of one curve valuation was known to specialists for many years. Computations can be found, e. g., in [13]. The Poincaré series of one divisorial valuation was computed in [12]. The Poincaré series of a collection of curve valuations was computed in [4] (see also [5] for a “simpler” proof). It was shown that (for a collection consisting of more than one valuation) this series coincides with the Alexander polynomial in several variables of the corresponding algebraic link. This implies that the Poincaré series determines the topology of the union of the corresponding curves.
@@
Here we consider Poincaré series of valuations on the algebra ${\mathcal{E}}_{\mathbb{K}^{2},0}$ of germs of holomorphic functions in two variables whose Taylor coefficients are from a subfield $\mathbb{K}$ of the field $\mathbb{C}$ of complex numbers (or, equivalently, on the algebra $\mathbb{K}[[x,y]]$ of formal power series). We compute the Poincaré series of one curve or one divisorial valuation on it: Theorems 1, 2, and 3. The Poincaré series are expressed in terms of a resolution of a special form and are represented as products/ratios of binomials of the form $(1-t^{m})$.
The computation of the Poincaré series of a curve valuation presented in [13] was combinatorial. However, it was known long before that. The importance of that paper was the observation that it coincided with the monodromy zeta function, i. e., with the Alexander polynomial of the corresponding algebraic knot divided by (1t)(1-t). All the computations of the Poincaré series made by the authors (starting from [4]) used some elements of topology: versions of the integration with respect to the Euler characteristic. The key property for that was the fact that the (additive) Euler characteristic of a complex affine space is equal to $1$. This does not work already for real analogues of the Poincaré series in [7] and [8] (the Euler characteristic of a real affine space is equal to $\pm 1$) and makes no (straightfoward) sense for affine spaces over subfields of $\mathbb{C}$. Therefore the computations in these cases (at least at the contemporary stage) are essentially combinatorial (modulo a certain analytic hint explained in Section 6).
The content of the paper is the following. In Section 2, we discuss curve valuations on the algebra ${\mathcal{E}}_{\mathbb{K}^{2},0}$. Section 3 is devoted to a description of a resolution process of an irreducible curve playing the central role in the computations. In Section 4, one associates Galois groups and subfields to the components of the exceptional divisor of a resolution. Section 5 is devoted to the computation of the semigroup Poincaré series, i. e., the sum $\sum_{v\in S^{\mathbb{K}}_{\nu}}t^{v}$ over the set $S^{\mathbb{K}}_{\nu}$ of values of the valuation $\nu$ on ${\mathcal{E}}_{\mathbb{K}^{2},0}$: the semigroup of values of the valuation. In Section 6, the classical Poincaré series (1) of a curve valuation is computed. In Section 7 it is computed for a divisorial valuation.
We are very thankful to the referee for reading carefully the paper and making a number of useful suggestions.
@@ 2 Curve valuations on functions over a field
Let $\mathbb{K}$ be a subfield of the field $\mathbb{C}$ of complex numbers and let ${\mathcal{E}}_{\mathbb{K}^{2},0}$ be the algebra of germs of holomorphic functions in two variables with the Taylor coefficients from $\mathbb{K}$. A (discrete, rank one) valuation on ${\mathcal{E}}_{\mathbb{K}^{2},0}$ (defined as above) is the restriction of a valuation on ${\mathcal{O}}_{\mathbb{C}^{2},0}$: see, e. g., [11]. In the framework of the discussion below, without changes it is possible to consider, instead of the algebra ${\mathcal{E}}_{\mathbb{K}^{2},0}$, the algebra $\mathbb{K}[[x,y]]$ of formal power series in two variables. We shall refer to ${\mathcal{E}}_{\mathbb{K}^{2},0}$ because of some convenience of the language.
Remark 1.
Strictly speaking, some constructions below (in particular the description of the action of the Galois group $G$ on branches at the beginning of section 3) cannot be applied directly to the algebra ${\mathcal{E}}_{\mathbb{K}^{2},0}$. The series for the branch $g\gamma$ and for the function $gh$ therein could be non-convergent. These constructions work in the algebra $K[[x,y]]$, however use of it everywhere produces some language problems. We do not pay attention to this problem since, for a bounded values of valuations we can treat not functions or series, but jets of order high enough, what means essentially polynomials. Informally one can say that the computations can be made in a sort of “an algebra of jets”.
Let $C$ be a plane algebroid branch on $(\mathbb{C}^{2},0)$, that is $C$ (after an appropriate change of the coordinates in $\mathbb{C}^{2}$ defined over $\mathbb{K}$) is given by
$x=x(\tau)=\tau^{m},\\ y=y(\tau)=\sum_{i\geq m}c_{i}\tau^{i}\in\mathbb{C}[[\tau]]\,$ (2)
with some $m>0$. (We assume that the parametrization (2) is primitive, i. e., the greatest common divisor of $m$ and of all $i$ such that $c_{i}\neq 0$ is equal to $1$.) The branch $C$ defines a valuation $\nu=\nu_{C}$ on the algebra ${\mathcal{O}}_{\mathbb{C}^{2},0}$ of holomorphic function germs in $x$ and $y$. For $f\in{\mathcal{O}}_{\mathbb{C}^{2},0}$, $\nu(f)$ is the degree of the leading term in the series $f(x(\tau),y(\tau))\in\mathbb{C}[[\tau]]$:
@@
The (classical) $\mathbb{K}$-Poincaré series of the branch $C$ is
$P_{C}^{\mathbb{K}}(t)=\sum_{v=0}^{\infty}\dim_{\mathbb{K}}\left(J^{\mathbb{K}}(v)/J^{\mathbb{K}}(v+1)\right)\cdot t^{v}.$ (4)
If $\mathbb{K}=\mathbb{R}$, $P^{\mathbb{R}}_{C}(t)$ is the classical Poincaré series of the valuation $\nu$ computed in [7].
@@
If $\mathbb{K}=\mathbb{C}$, all the coefficients of the series (4) are equal to 0 or 11. This is not the case, in general, for a proper subfield $\mathbb{K}\subset\mathbb{C}$. In particular, if $\mathbb{K}=\mathbb{R}$ (the field of real numbers), all the coefficients are equal to 0, 11, or 22 ([7]).
Let $S^{\mathbb{K}}_{C}:=\{\nu_{C}(f):f\in{\mathcal{E}}_{\mathbb{K}^{2},0}\}\subset\mathbb{Z}_{\geq 0}$ be the semigroup of values of the valuation $\nu_{C}$ on ${\mathcal{E}}_{\mathbb{K}^{2},0}$.
@@
(cf. [7, Example 1 and Remark 3]) Let $\mathbb{K}=\mathbb{R}$, and let the branch $C$ be given by the parametrization $(\tau^{4},\alpha\tau^{6}+\tau^{7})$ with a generic complex $\alpha$. The semigroup of values of the valuation $\nu_{C}$ on ${\mathcal{E}}_{\mathbb{R}^{2},0}$ is equal to $S_{C}^{\mathbb{R}}=\langle 4,6,25\rangle$. Here $\langle\bullet,\bullet,\ldots\rangle$ denotes the subsemigroup of $\mathbb{Z}_{\geq 0}$ generated by the corresponding natural numbers:
$\langle m_{1},\ldots,m_{p}\rangle=\{m\in\mathbb{Z}_{\geq 0}:m=k_{1}m_{1}+\ldots+k_{p}m_{p}\text{\ \ with\ \ }k_{i}\in\mathbb{Z}_{\geq 0}\}.$
This semigroup coincides with the semigroup of values of the curve valuation $\nu_{C^{\prime}}$ on ${\mathcal{O}}_{\mathbb{C}^{2},0}$, where the curve $C^{\prime}$ is given by the parametrization $(\tau^{4},\tau^{6}+\tau^{19})$. A direct relation between the curves $C$ and $C^{\prime}$ is not clear. The semigroup Poincaré series (both of the valuation $\nu_{C}$ on ${\mathcal{E}}_{\mathbb{R}^{2},0}$ and of the valuation $\nu_{C^{\prime}}$ on ${\mathcal{O}}_{\mathbb{C}^{2},0}$) is given by
$P^{S}_{C}(t)=\frac{(1-t^{12})(1-t^{50})}{(1-t^{4})(1-t^{6})(1-t^{25})}\,.$
The classical Poincaré series of the valuation $\nu_{C}$ on ${\mathcal{E}}_{\mathbb{R}^{2},0}$ is given by
$P^{\mathbb{R}}_{C}(t)=\frac{(1-t^{24})(1-t^{50})}{(1-t^{4})(1-t^{6})(1-t^{25})}\,.$
@@ 3 The resolution process of a branch
Let $G$ be the Galois group of the extension $\mathbb{C}/\mathbb{K}$. i. e., the group of automorphisms of the field $\mathbb{C}$ which are trivial on $\mathbb{K}$. The group $G$ acts on branches (irreducible complex plane curve germs). This action can be defined by any of the following three ways.
@@
    Let a branch $\gamma$ be given by a parametrization $x=\tau^{k}$, $y=\sum\limits_{i\geq k}a_{i}\tau^{i}$. Then the branch $g\gamma$, $g\in G$, is given by the parametrization $x=\tau^{k}$, $y=\sum\limits_{i\geq k}g(a_{i})\tau^{i}$.
@@
    Let a branch $\gamma$ be given by an equation $h(x,y)=0$, where $h(x,y)=\sum\limits_{i,j\geq 0}b_{ij}x^{i}y^{j}$. Then the branch $g\gamma$ is given by the equation $gh(x,y)=0$, where $gh(x,y):=\sum\limits_{i,j\geq 0}g(b_{ij})x^{i}y^{j}$.
@@
    Let a branch $\gamma$ be considered as a subset of $(\mathbb{C}^{2},0)$: $(\gamma,0)\subset(\mathbb{C}^{2},0)$. The group $G$ acts on the plane $\mathbb{C}^{2}$ by $g(x,y)=(g(x),g(y))$. Then the branch $g\gamma$ as a subset of $(\mathbb{C}^{2},0)$ is just the image of $\gamma$ under this action.
Remark 3.
For fixed $x$ and $y$ the parametrization of $\gamma$ is not unique. One can obtain another parametrization from one of them multiplying the parameter $\tau$ by a root $\xi$ of degree $k$ of $1$. For $g\in G$, $g(\xi)$ is also a root of degree $k$ of $1$. Thus this only leads to another parametrization the branch $g\gamma$. It is not difficult to see that, for a $G$-invariant curve germ $\mathcal{C}$} (possibly reducible), there exists a function germ $h\in{\mathcal{E}}_{\mathbb{K}^{2},0}$ such that ${\mathcal{C}}=\{h=0\}$.
Let us consider the minimal $G$-invariant resolution process of the branch $C$ defining the valuation (by a sequence of blow-ups). This means the following. When we blow-up the origin in $\mathbb{C}^{2}$, we have got one of the following three possibilities.
  1. 1)
    There are finitely many different branches in the $G$-orbit of the branch $C$, their strict transforms are smooth and intersect the exceptional divisor transversally at pairwise different points. We get the (minimal embedded) $G$-invariant resolution of $C$ and the process is completed.
  2. 2)
    The strict transforms of the branches from the $G$-orbit of $C$ (whose number may be infinite) intersect the exceptional divisor at finite number of points, but either they are not smooth, or do not intersect the exceptional divisor transversally, or some (different) of them intersect it at the same point. (In this case, if the $G$-orbit of $C$ has infinitely many branches, the number of strict transforms intersecting the exceptional divisor at one point is also infinite.) In this case one has to continue the resolution process.
  3. 3)
    The strict transforms of the branches from the $G$-orbit of C{C} intersect the exceptional divisor at infinitely many different points. In this case the process is interrupted. (Do not mix with the term “completed”. In this case, one cannot get a $G$-invariant resolution since blowing-ups infinitely many points does not produce a complex surface.)
In Case 2) one has to continue the resolution process blowing-up (all) the intersection points of the strict transforms of the branches from the $G$-orbit of $C$ with the exceptional divisor (their number being finite). One gets one of the analogues of the cases 1)–3) again. Etc.
As the result, we arrive to one of the following three situations.
  1. I.
    There are finitely many branches in the $G$-orbit of $C$ and they are resolved at a certain step.
  2. II.
    The $G$-orbit of $C$ has infinitely many branches, but the described process goes on without interruption. This means that at each step one has to blow-up finitely many points, but there are infinitely many steps like that. One does not arrive to a $G$-invariant resolution of $C$, but has got a $G$-invariant resolution process (an infinite one).
  3. III.
    The process is interrupted at a certain step. This means that at this step the strict transforms of the branches from the $G$-orbit of $C$ intersect each of the new-born components of the exceptional divisor at infinitely many points. In this case one gets neither a $G$-invariant resolution of $C$, nor a resolution process.
The cases I–III take place in the following situations.
  1. I.
    All the coefficients in the series (2) are from a finite extension of $\mathbb{K}$. This is equivalent to the condition that the $G$-orbit of $C$ has finitely many elements.
  2. II.
  3. III.
At each step of a $G$-invariant resolution of $C$, the action of the Galois group $G$ on $\mathbb{C}^{2}$ lifts to its action on the surface of modification and thus induces an action on the set of irreducible components of the exceptional divisor. These components are in bijection with the vertices of the dual graph of the modification and thus one has an action of $G$ on this dual graph.
In the case I, the dual graph $\Gamma$ of the resolution looks like in Figure 1.
@@
For a description of the resolution (or of the resolution process), it is convenient to use the quotient $\check{\Gamma}$ of the graph $\Gamma$ by the action of the Galois group $G$. In this case it looks like in Figure 2.
@@
The graph $\Gamma$ contains a subgraph which geometrically coincides with the dual graph of a usual resolution of the curve $C$: the lower part of $\Gamma$ on Figure 1. (We do not consider this resolution below.) We use “short” notations (by greek letters) for vertices on this subgraph (and also for the corresponding vertices of the graph $\check{\Gamma}$). All other vertices are obtained from them by the action of elements of the Galois group $G$.
@@
In Figures 1 and 2, $\sigma_{i}$, $i=0,1,\ldots,g$, (and also their images under the action of $G$) are the dead ends of the graph ($g$ is the number of the Puiseux pairs of the curve $C$), $\tau_{i}$, $i=1,\ldots,g$, (and also their images) are the rupture points of it, $\rho_{j}$, $j=1,\ldots,s$, are the splitting points corresponding to the changes of the isotropy subgroups for the $G$-action on $\Gamma$ (that is the splitting points for usual resolutions of different branches from the orbit of $C$, there are finitely many of them in Case I). Let $\delta_{C}$ be the vertex corresponding to the component of the exceptional divisor intersecting the strict transform of the curve $C$. In Figure 1, $\delta_{C}=\tau_{6}$. Figure 2 shows $\check{\Gamma}$ in the case when the process of the $G$-resolution of $C$ is not finished when $C$ itself is resolved, $\delta_{C}=\rho_{s}$ there. (Otherwise $\delta_{C}$ coincides with $\tau_{g}$.)
@@
Figure 2: The quotient $\check{\Gamma}$ of the resolution graph in Case I (with arbitrary $g$ and $s$).
@@
In Case II, one has the dual graph of the resolution process. The corresponding quotient $\check{\Gamma}$ looks like in Figure 3. One can say that $s$ is equal to infinity in this case.
@@
Figure 3: The quotient $\check{\Gamma}$ of the graph of the resolution process in Case II (with $s=\infty$).
In Case III, the quotient of the modification graph looks like in Figure 4. The curve $C$ may be not resolved at the interruption step.
@@
In the latter case, the valuation coincides with $n\nu$, where $\nu=\nu_{\rho_{s}}$ is the divisorial valuation defined by $E_{\rho_{s}}$ and $n$ is the intersection multiplicity $E_{\rho_{s}}\circ\widetilde{C}$ of the strict transform $\widetilde{C}$ of the curve $C$ with the component $E_{\rho_{s}}$.
Figure 4: The quotient $\check{\Gamma}$ of the modification graph in Case III.
Example 2.
Let $\mathbb{K}=\mathbb{Q}$ and the branch $C$ given by the parametrization $x=\tau^{4}$, $y=\alpha_{0}\tau^{6}+t^{7}+\sum_{i\geq 1}\alpha_{i}\tau^{7+i}$ where $\alpha_{0},\alpha_{1},\ldots\in\mathbb{C}\setminus\{0\}$. Let $\mathbb{K}^{\prime}:=\mathbb{Q}(\alpha_{0},\alpha_{1},\ldots)\subset\mathbb{C}$.
Figure 5: The dual graph of $C$
One can distinguish the following cases:
  1. 1.
    $\mathbb{K}^{\prime}$ is finite algebraic extension of $\mathbb{Q}$. In this case $\mathbb{K}^{\prime}=\mathbb{Q}(\alpha_{0},\,\ldots,\alpha_{k})$ for some $k$ and if we add the conditions $\alpha_{0}\notin\mathbb{Q}$ (algebraic, generic) and $\alpha_{i}\notin\mathbb{Q}(\alpha_{0},\,\ldots,\alpha_{i-1})$ (algebraic, generic) for $1\leq i\leq k$ then $s=k+1$ and the splitting points are $\rho_{1}=3$, $\rho_{i}=5+(i-1)$ for $i=2,\ldots,k+1$. Thus we are in the Case I.
    Pay attention that, if $\alpha_{0}\in\mathbb{K}^{\prime}$ is not generic, the pictire can be somewaht different. For example, let $y=\sqrt{2}\tau^{6}+\tau^{7}$. Then the branch $C$ has the equation $f(x,y)=(y^{2}-2x^{3})^{2}-4\sqrt{2}yx^{5}-x^{7}=0$ and the conjugate of $C$ with respect to $G$, gCgC ($g\in G$ is given by $g(\sqrt{2})=-\sqrt{2}$), has the parametrization $(\tau^{4},-\sqrt{2}\tau^{6}+\tau^{7})$ and the equation $f^{\prime}(x,y)=gf(x,y)=(y^{2}-2x^{3})^{2}+4\sqrt{2}yx^{5}-x^{7}=0$. In this case the only splitting point is $\rho=5$ (also $\delta_{C}=5$). Notice that the intersection multiplicity of $C$ and $C^{\prime}=gC$ is $26$ and that the coordinate of the intersection point of the strict transform of $C$ (resp. $C^{\prime}$) is a certain $a\in\mathbb{Q}(\sqrt{2})\setminus\mathbb{Q}$ (resp. $g(a)$).
  2. 2.
    $\mathbb{K}^{\prime}$ is an infinite algebraic extension of $\mathbb{Q}=\mathbb{K}$. Then $s=\infty$ and we are in Case II. As in the previous case, in general, the splitting points are $3,6,7,\ldots$.
  3. 3.
    $\mathbb{K}^{\prime}$ is a trascendental extension of $\mathbb{Q}$. In this case we are in Case III. For simplicity, let us assume that $\alpha_{0}$ is trascendental over $\mathbb{Q}$. Then the process is interrumpted in the divisor $\delta_{\nu}=3$. In order to get a function $g\in{\mathcal{O}}_{\mathbb{C}^{2},0}$ such that $\nu(g)=2r+1\geq 7$ one needs that
    $g=(y^{2}-\alpha_{0}^{2}x^{3})+\text{ higher order terms }\;.$
    But in this case the equation can not be rational. So, in this case, the semigroup $S^{\mathbb{Q}}_{\nu}$ is $\langle 4,6\rangle\subset 2\mathbb{Z}$ and the valuation $\nu$ is $2$ times the divisorial valuation defined by the component $E_{3}$.
@@
Let $F\in\mathbb{K}[[x,y]]$ be irreducible. The local ring of the irreducible curve (over $\mathbb{K}$) defined by $F$ is $R=\mathbb{K}[[x,y]]/(F)$. The integral closure $\overline{R}$ in its quotient field is the valuation ring $R_{\nu}$ of a discrete valuation $\nu$, its residual field $\mathbb{K}^{\prime}$ is a finite algebraic extension of $\mathbb{K}$ and could be realized and therefore identified with a subfield of $\mathbb{C}$. The equation $F=0$ defines a reduced algebroid curve over the complex field $\mathbb{C}$, each branch corresponds to an irreducible factor of $F$ in $\mathbb{C}[[x,y]]$. Let $C$ be one of these branches. The valuation $\nu_{C}$ defined by $C$ on $\mathbb{K}[[x,y]]$ coincides with the valuation $\nu$ (seen as a valuation on $\mathbb{K}[[x,y]]$ forgetting the quotient by $F$) and the Poincaré series (over $\mathbb{K}$) defined by $\nu_{C}$ is the Poincaré series defined by $\nu$. Note that all the branches are equisingular, the minimal resolution of $\{F=0\}$ (in the complex setting) coincides with the $G$-resolution of $\nu_{C}$ and all the valuations $\nu_{C}$, for complex branches $C$ of $\{F=0\}$, restrict to the same valuation $\omega$ in $\mathbb{K}[[x,y]]$. Moreover, if $G^{\prime}$ is the Galois group of the extension $\mathbb{K}\subset{\mathbb{L}}$, $\mathbb{L}$} the normal closure of $\mathbb{K}^{\prime}$ in $\mathbb{C}$, $G^{\prime}$ is the quotient of $G$ by the isotropy group of the field $\mathbb{L}$. This situation just corresponds to the case I described above.
Let $C$ be a curve parametrized as in (2). The map $f\mapsto f(x(\tau),y(\tau))$ defines an homomorphism $\Phi:\mathbb{K}[[x,y]]\to\mathbb{C}[[\tau]]$. If $\ker(\Phi)\neq 0$, then one has $\ker(\Phi)=(F)$ for some irreducible $F\in\mathbb{K}[[x,y]]$ and we are in the above described situation, i. e., also in Case I. Otherwise, $\ker(\Phi)=(0)$ and we are in Case II or III. In this case the valuation $\nu_{C}$ is centered in the two-dimensional ring $\mathbb{K}[[x,y]]$. Case II corresponds to the case in which the residual field $\mathbb{K}_{\nu_{C}}$ of the valuation ring is an algebraic extension of $\mathbb{K}$ but not finite dimensional. Case III corresponds to the case in which $\mathbb{K}_{\nu_{C}}$ is a trascendental extension of $\mathbb{K}$.